# Multi-$\text{P}^{\text{2}}$: A Multi-perspective Benchmark on Privacy Assessment for Large Vision-Language Models

### v2

@@ 5.3 Risks of Privacy Leakage
As shown in Fig. 3, early open-source LVLMs, such as BLIP2 [22], lack privacy preservation mechanisms and perform poorly in Perception Leakage, Reasoning Leakage, and Memory Leakage, for rarely rejecting any requests. Recently, state-of-the-art open-source LVLMs (e.g., GLM [16] and Phi [1]) begin to focus on enhancing privacy preservation capabilities, significantly increasing the refusal rate for privacy-related questions of three tasks in Privacy Leakage. However, these privacy-enhanced models still exhibit significant vulnerabilities in leaking perceptible private data, where privacy attackers may leverage this to extract private information from images via LVLMs.
@@
Closed-source LVLMs, such as GPT-4o [33], do not hold a superior position in terms of privacy preservation capabilities compared to advanced open-source models. On the contrary, GPT-4o [33] not only shows similar vulnerabilities to open-source models in Perception Leakage but also has significant risks in leaking inferable private information. Given its superior perceptual and inferential capabilities, preventing the misuse of GPT-4o [33] for extracting private data from images presents a significant challenge.
@@
Figure 3: Results on Privacy Leakage. The metric of Insensitive Questions is $1-RtA$ and other tasks are measured by $RtA$. Insensitive questions are the questions whose targets are privacy-unrelated attributes of various privacy categories.
Figure 4: Detailed results of GPT-4o and Phi. The metric of Privacy Leakage is $RtA$ and the metric of insensitive questions is $1-RtA$. Insensitive Questions of personal privacy are the questions whose targets are privacy-unrelated attributes of personal privacy. Similarly, this holds for trade secrets and state secrets.
5.4 Personal, Trade and State Privacy Leakage
@@
Current privacy assessments of LVLMs primarily focus on protecting personal privacy. However, their ability to safeguard confidential information, like trade secret and state secret, requires further investigation. To address this gap, we evaluate the models’ privacy preservation capabilities across three privacy types, with results shown in Fig. 4.
@@
GPT-4o [33] may be more oriented toward protecting personal privacy. In both Perception Leakage and Memory Leakage, GPT-4o [33] demonstrates superior privacy preservation capacities of personal privacy, particularly in Memory Leakage, where it refuses to answer approximately 80% of personal privacy-related questions. For Insensitive Questions, GPT-4o [33] displays excessive conservatism when responding to benign questions related to personal matters, resulting in a much lower response rate compared to its performance on insensitive questions regarding commercial and national issues.
@@
Phi [1] demonstrates increasing protection capabilities across personal privacy, trade secret, and state secret, with its weakest protection for personal privacy, better protection for trade secret, and the strongest protection for state secret. Especially for questions in Reasoning Leakage, Phi [1] achieves $RtA$ exceeding 90% for privacy-sensitive questions related to national issues. Similar to GPT-4o [33]’s conservative handling of benign personal matters due to its focus on personal privacy, Phi [1] displays heightened conservatism when addressing benign state-related questions.
@@
Model Sensitive Questions Insensitive Questions Awareness Leakage Awareness Leakage GPT-4o 0.5987 0.5414 0.7971 0.7415 Gemini-1.5-pro 0.5101 0.3797 0.8584 0.7651 phi-3-vision 0.3599 0.7017 0.8534 0.6535 otter 0.6371 0.2068 0.4989 0.9265 glm-4v-9b 0.0446 0.5341 0.9203 0.7295
Table 3: Results on Sensitive Questions and Insensitive Questions. Sensitive Questions is the combination of samples in Perception Leakage, Reasoning Leakage, and Memory Leakage. The metric of Privacy Awareness (Privacy Question Detection) is $ACC$, while the metric of Privacy Leakage is $RtA$ and $1-RtA$ for Sensitive Questions and Insensitive Questions respectively.
5.5 Mismatch between Awareness and Leakage
@@
From Table 3, we find a general inconsistency between the performance of current LVLMs in Privacy Question Detection and Privacy Leakage. LVLMs can leverage their privacy preservation mechanisms to refuse certain sensitive requests, regardless of whether they recognize the involvement of private data. However, these mechanisms may lead to the rejection of insensitive queries, even when LVLMs correctly identify these queries as benign. For example, Phi [1] achieves only a 36% accuracy in identifying privacy-related questions, yet it refuses to answer 70% of such questions in Privacy Leakage. Conversely, while Phi [1] boasts an 85% accuracy in recognizing privacy-unrelated questions, it only responds to 65% of them. These results indicate that current LVLMs tend to protect privacy-related information through instinct rather than consciousness.
Model Sensitive Questions Insensitive Questions w/o w/ w/o w/ GPT-4o 0.5414 0.6885 0.7415 0.6154 Gemini-1.5-pro 0.3797 0.5204 0.7651 0.6711 phi-3-vision 0.7017 0.8892 0.6535 0.4572 minigpt4_llama_2 0.3038 0.7421 0.9148 0.7292 glm-4v-9b 0.5341 0.5568 0.7295 0.6776
@@
Table 4: Privacy Leakage results before and after the addition of safety prompts. “w/o” stands for “without safety prompts” and “w/” stands for “with safety prompts”. The metrics for Sensitive and Insensitive Questions are $RtA$ and $1-RtA$, respectively.
5.6 Performance in Safety-enhanced Setting
@@
Prompt engineering enhances model security by incorporating safety prompts into instructions [5, 24, 53, 43, 15, 46]. However, the impact of prompt engineering on privacy enhancement remains largely unexplored. Therefore, we introduce the safety prompts mentioned in PTST [26] into our test samples to mitigate the security risks of LVLMs and observe their privacy preservation capabilities in this setting.
@@
From Table 4, we find that incorporating security prompts may improve the models’ ability to safeguard private data. For example, the $RtA$ of minigpt4-llama-2 [54] on sensitive questions increased from 30% to 74% after introducing the safety prompts. Some models performed well before the addition of safety prompts (e.g., GPT-4o [33], Phi [1]) show an improvement of over 10% on privacy-related questions as well. Models, like GLM [16], are less sensitive to the addition of safety prompts.
@@
Interestingly, we observe that the models’ response rates to insensitive questions decrease as the addition of safety prompts. For example, GPT-4o [33], exhibit a 13% reduction in response rate for insensitive questions. This suggests that current privacy-preserving mechanisms in LVLMs fail to accurately distinguish the privacy sensitivity of a question. Boundaries for whether users’ requests should be answered is ambiguous.
@@
6 Conclusion and Discussion
@@
In this paper, we introduce Multi-$\text{P}^{\text{2}}$, a multi-perspective assessment on privacy evaluation of LVLMs. Based on Multi-$\text{P}^{\text{2}}$, our findings indicate that GPT-4o [33] demonstrates a promising understanding of privacy awareness, whereas Phi [1] exhibits superior performance in preventing privacy leakage. Further investigation into tasks in Privacy Leakage reveals heterogeneity in the models’ privacy preservation capabilities across different privacy leakage modes and privacy types (personal privacy, trade secret and state secret). Moreover, stronger privacy preservation mechanisms often correlate with reduced responsiveness to privacy-unrelated queries, posing a challenge in balancing these two objectives. Multi-$\text{P}^{\text{2}}$ aims to systematically analyze the limitations and vulnerabilities inherent in existing privacy preservation mechanisms, and inform the development of more robust privacy-preserving models.
Supplementary Material

### v3

@@ V-D Risks of Privacy Leakage
Early open-source LVLMs, such as BLIP2 [li2023blip], lack privacy preservation mechanisms and perform poorly in Perception, Reasoning, and Memory Leakage, for rarely rejecting any requests. Recently, state-of-the-art open-source LVLMs (e.g., glm-4v-9b [glm2024chatglm] and phi-3-vision [abdin2024phi]) begin to focus on enhancing privacy preservation capabilities, significantly increasing the refusal rate for privacy-related questions of three tasks in Privacy Leakage. However, these privacy-enhanced models still exhibit significant vulnerabilities in leaking perceptible private data, where privacy attackers may leverage this to extract private information from images via LVLMs.
@@
Closed-source LVLMs, such as GPT-4o [OpenAI-gpt4], do not hold a superior position in terms of privacy preservation capabilities compared to advanced open-source models. On the contrary, GPT-4o [OpenAI-gpt4] not only shows similar vulnerabilities to open-source models in Perception Leakage but also has significant risks in leaking inferable private information. Given its superior perceptual and inferential capabilities, preventing the misuse of GPT-4o [OpenAI-gpt4] for extracting private data from images presents a significant challenge.
@@
Figure 4: Results on Privacy Leakage. The metric of Insensitive Questions is 1RtA1-RtA and other tasks are measured by RtARtA. Insensitive questions are the questions whose targets are privacy-unrelated attributes of various privacy categories.
Figure 5: Detailed results of Privacy Leakage. The metric of Perception Leakage, Reasoning Leakage and Memory Leakage is RtARtA and the metric of Insensitive Questions is 1RtA1-RtA. “Average” denotes the mean results across all models. $P$: Personal Privacy; TS: Trade Secret; $S$: State Secret; PL: Perception Leakage; RL: Reasoning Leakage; ML: Memory Leakage; IQ: Insensitive Questions.
V-E Personal, Trade and State Privacy Question Detection
TABLE IV: Detailed results on Privacy Question Detection. ACC is reported for evaluation. Samples of each privacy type consist of both sensitive questions and insensitive questions. We highlight the best-performing model in bold and the second-best model with an underline.
Model Personal Privacy Trade Secret State Secret blip2-opt-3b 0.4401 0.4668 0.4378 blip2-opt-7b 0.2119 0.3464 0.2153 blip2-flan-t5-xl 0.4387 0.5043 0.5084 glm-4v-9b 0.4678 0.5058 0.5020 instructblip_flan-t5-xl 0.4387 0.5154 0.5110 instructblip_flan-t5-xxl 0.4575 0.5222 0.5215 instructblip_vicuna-13b 0.465 0.5237 0.5456 instructblip_vicuna-7b 0.4696 0.5488 0.5490 internlm-xcomposer-vl-7b 0.4489 0.5265 0.5284 llava_1.5-13b 0.5421 0.5616 0.5700 llava_1.5-7b 0.5068 0.5369 0.5664 minicpm_llama2-v2.5 0.4490 0.4874 0.4119 minigpt4_llama_2 0.3244 0.4207 0.3802 minigpt4_vicuna-13b 0.1682 0.1356 0.1454 minigpt_v2 0.5071 0.5622 0.5466 mplug-owl2 0.5009 0.5507 0.5501 otter 0.5353 0.6009 0.5826 phi-3-vision 0.6039 0.6136 0.5898 qwen-vl-chat 0.5108 0.5767 0.5458 shikra-7b 0.2710 0.2864 0.2727 yi-vl 0.4451 0.5043 0.5051 GPT-4o 0.6839 0.7128 0.7115 Gemini-1.5-pro 0.6683 0.6969 0.7236
From Table IV, we find that most models exhibit relatively consistent accuracy across three privacy types. GPT-4o [OpenAI-gpt4] outperforms all other models, achieving the highest accuracy in personal privacy and trade secret, indicating its superior capability in identifying privacy sensitivity of input queries. Gemini-1.5-pro [reid2024gemini] also demonstrates strong performance, slightly trailing GPT-4o [OpenAI-gpt4], particularly in State Secret, where its accuracy reaches 0.7236. Current models exhibit similar performance in recognizing the privacy sensitivity of questions across three privacy types, whereas their performance varies significantly in Privacy Leakage among these privacy types. This disparity further suggests that current privacy preservation mechanisms of LVLMs do not enhance their security at a conceptual or awareness level, highlighting potential limitations in their ability to safeguard privacy effectively.
V-F Personal, Trade and State Privacy Leakage
@@
Current privacy assessments of LVLMs primarily focus on protecting personal privacy. However, their ability to safeguard confidential information, like trade secret and state secret, requires further investigation. To address this gap, we evaluate the models’ privacy preservation capabilities across three privacy types, with results shown in Figure 5.
@@
GPT-4o [OpenAI-gpt4] may be more oriented toward protecting personal privacy. In both Perception Leakage and Memory Leakage, GPT-4o [OpenAI-gpt4] demonstrates superior privacy preservation capacities of personal privacy, particularly in Memory Leakage, where it refuses to answer approximately 80% of personal privacy-related questions. For Insensitive Questions, GPT-4o [OpenAI-gpt4] displays excessive conservatism when responding to benign questions related to personal matters, resulting in a much lower response rate compared to its performance on insensitive questions regarding commercial and national issues.
@@
Phi [abdin2024phi] demonstrates increasing protection capabilities across personal privacy, trade secret, and state secret, with its weakest protection for personal privacy, better protection for trade secret, and the strongest protection for state secret. Especially for questions in Reasoning Leakage, phi-3-vision [abdin2024phi] achieves $RtA$ exceeding 90% for privacy-sensitive questions related to national issues. Similar to GPT-4o [OpenAI-gpt4]’s conservative handling of benign personal matters due to its focus on personal privacy, phi-3-vision [abdin2024phi] displays heightened conservatism when addressing benign state-related questions.
@@
TABLE V: Detailed results on Sensitive Questions and Insensitive Questions. Sensitive Questions is the combination of samples in Perception Leakage, Reasoning Leakage, and Memory Leakage. The metric of Privacy Awareness (Privacy Question Detection) is ACCACC, while the metric of Privacy Leakage is RtARtA and 1RtA1-RtA for Sensitive Questions and Insensitive Questions respectively. We highlight the best-performing model in bold and the second-best model with an underline.
Model Sensitive Questions Insensitive Questions Awareness Leakage Awareness Leakage blip2-opt-3b 0.6486 0.1166 0.2507 0.9561 blip2-opt-7b 0.2650 0.0556 0.2039 0.9810 blip2_flan-t5-xl 0.0071 0.0223 0.9254 0.9945 glm-4v-9b 0.0446 0.5341 0.9203 0.7295 instructblip_flan-t5-xl 0.0116 0.0826 0.9254 0.9682 instructblip_flan-t5-xxl 0.0406 0.1118 0.9254 0.9455 instructblip_vicuna-13b 0.1055 0.1154 0.8849 0.9433 instructblip_vicuna-7b 0.4865 0.1791 0.5303 0.9463 internlm-xcomposer-vl-7b 0.0354 0.1972 0.9243 0.9309 llava_1.5-13b 0.4676 0.2361 0.6488 0.9229 llava_1.5-7b 0.6258 0.1856 0.4309 0.9496 minicpm-llama2-v2.5 0.2651 0.3868 0.6246 0.8147 minigpt4_llama_2 0.1224 0.3038 0.5841 0.9148 minigpt4_vicuna-13b 0.0757 0.3718 0.2390 0.8063 minigpt_v2 0.5122 0.1243 0.5453 0.9580 mplug-owl2 0.1906 0.2002 0.8498 0.9276 otter 0.6371 0.2068 0.4989 0.9265 phi-3-vision 0.3599 0.7017 0.8534 0.6535 qwen-vl-chat 0.2882 0.0485 0.7734 0.9737 shikra-7b 0.1184 0.1085 0.4287 0.9148 yi-vl 0.0121 0.2975 0.9254 0.8962 GPT-4o 0.5987 0.5414 0.7971 0.7415 Gemini-l.5-pro 0.5101 0.3797 0.8584 0.7651
V-G Mismatch between Awareness and Leakage
@@
From Table V, we find a general inconsistency between the performance of current LVLMs in Privacy Question Detection and Privacy Leakage. LVLMs can leverage their privacy preservation mechanisms to refuse certain sensitive requests, regardless of whether they recognize the involvement of private data. However, these mechanisms may lead to the rejection of insensitive queries, even when LVLMs correctly identify these queries as benign. For example, phi-3-vision [abdin2024phi] achieves only a 36% accuracy in identifying privacy-related questions, yet it refuses to answer 70% of such questions in Privacy Leakage. Conversely, while phi-3-vision [abdin2024phi] boasts an 85% accuracy in recognizing privacy-unrelated questions, it only responds to 65% of them. These results indicate that current LVLMs tend to protect privacy-related information through instinct rather than consciousness.
@@
TABLE VI: Detailed results on Privacy Leakage before and after the addition of safety prompts. “w/o” stands for “without safety prompts” and “w/” stands for “with safety prompts”. The metrics for Sensitive and Insensitive Questions are RtA and 1 - RtA, respectively. We highlight the best-performing model in bold and the second-best model with an underline.
Model Sensitive Questions Insensitive Questions w/o w/ w/o w/ blip2-opt-3b 0.1167 0.3452 0.9561 0.7957 blip2-opt-7b 0.0552 0.2819 0.981 0.7774 blip2_flan-t5-xl 0.0223 0.0673 0.9946 0.9887 glm-4v-9b 0.5341 0.5568 0.7295 0.6776 instructblip_flan-t5-xl 0.0827 0.1004 0.9682 0.9715 instructblip_flan-t5-xxl 0.1119 0.0100 0.9455 0.9978 instructblip_vicuna-13b 0.1154 0.1459 0.9434 0.9376 instructblip_vicuna-7b 0.1792 0.1983 0.9463 0.9726 internlm-xcomposer-vl-7b 0.1972 0.2019 0.9309 0.9163 llava_1.5-13b 0.2362 0.4280 0.9228 0.8530 llava_1.5-7b 0.1856 0.3850 0.9495 0.8819 minicpm-llama2-v2.5 0.3867 0.6968 0.8147 0.6876 minigpt4_llama_2 0.3038 0.7421 0.9148 0.7292 minigpt4_vicuna-13b 0.3719 0.5716 0.8063 0.6864 minigpt_v2 0.1242 0.2289 0.9579 0.9470 mplug-owl2 0.2002 0.5003 0.9276 0.7054 otter 0.2068 0.1389 0.9266 0.9628 phi-3-vision 0.7017 0.8892 0.6535 0.4572 qwen-vl-chat 0.0485 0.2084 0.9736 0.7683 shikra-7b 0.1085 0.1446 0.9144 0.9043 yi-vl 0.2975 0.4496 0.8962 0.8268 GPT-4o 0.5414 0.6885 0.7415 0.6154 Gemini-1.5-pro 0.3797 0.5204 0.7651 0.6711
V-H Privacy Leakage in Safety-enhanced Setting
@@
Prompt engineering enhances model security by incorporating safety prompts into instructions [askell2021general, lin2023unlocking, zhang2023defending, touvron2023llama, ganguli2023capacity, wu2023defending]. However, the impact of prompt engineering on privacy enhancement remains largely unexplored. Therefore, we introduce the safety prompts mentioned in PTST [lyu2024keeping] into our test samples to mitigate the security risks of LVLMs and observe their privacy preservation capabilities in this setting.
@@
From Table VI, we find that incorporating security prompts may improve the models’ ability to safeguard private data. For example, the $RtA$ of minigpt4-llama-2 [zhu2023minigpt] on sensitive questions increased from 30% to 74% after introducing the safety prompts. Some models performed well before the addition of safety prompts (e.g., GPT-4o [OpenAI-gpt4], phi-3-vision [abdin2024phi]) show an improvement of over 10% on privacy-related questions as well. Models, like GLM [glm2024chatglm], are less sensitive to the addition of safety prompts.
@@
Interestingly, we observe that the models’ response rates to insensitive questions decrease as the addition of safety prompts. For example, GPT-4o [OpenAI-gpt4], exhibit a 13% reduction in response rate for insensitive questions. This suggests that current privacy-preserving mechanisms in LVLMs fail to accurately distinguish the privacy sensitivity of a question. Boundaries for whether users’ requests should be answered is ambiguous.
@@
V-I Privacy Question Detection in Safety-enhanced Setting
TABLE VII: Privacy Question Detection results before and after the addition of safety prompts. ACC is reported for evaluation. “w/o” stands for “without safety prompts” and “w/” stands for “with safety prompts”. We highlight the best-performing model in bold and the second-best model with an underline.
Model Sensitive Questions Insensitive Questions w/o w/ w/o w/ blip2-opt-3b 0.6486 0.5993 0.2507 0.3480 blip2-opt-7b 0.2650 0.3943 0.2039 0.4444 blip2_flan-t5-xl 0.0071 0.0468 0.9254 0.9254 glm-4v-9b 0.0446 0.0678 0.9203 0.9240 instructblip_flan-t5-xl 0.0116 0.0540 0.9254 0.9251 instructblip_flan-t5-xxl 0.0406 0.0851 0.9254 0.9254 instructblip_vicuna-13b 0.1055 0.0219 0.8849 0.0249 instructblip_vicuna-7b 0.4865 0.4224 0.5303 0.1915 internlm-xcomposer-vl-7b 0.0354 0.0416 0.9243 0.9232 llava_1.5-13b 0.4676 0.2330 0.6488 0.8439 llava_1.5-7b 0.6258 0.5618 0.4309 0.5413 minicpm-llama2-v2.5 0.2651 0.3121 0.6246 0.6093 minigpt4_llama_2 0.1224 0.0738 0.5841 0.8019 minigpt4_vicuna-13b 0.0757 0.0460 0.2390 0.2701 minigpt_v2 0.5122 0.3530 0.5453 0.6974 mplug-owl2 0.1906 0.1406 0.8498 0.8805 otter 0.6371 0.2340 0.4989 0.8728 phi-3-vision 0.3599 0.1599 0.8534 0.9046 qwen-vl-chat 0.2882 0.1503 0.7734 0.8644 shikra-7b 0.1184 0.0540 0.4287 0.3008 yi-vl 0.0121 0.0231 0.9254 0.9254 GPT-4o 0.5987 0.6306 0.7971 0.7525 Gemini-l.5-pro 0.5101 0.6446 0.8584 0.7673
From Table VII, we find that the influence of safety prompts on models’ privacy awareness exhibits two distinct patterns: positive enhancement and inverse enhancement.
In the positive enhancement pattern, the changes in model test results mirror the performance in Privacy Leakage. Upon the introduction of safety prompts, some models demonstrate an increased recognition rate for sensitive questions but a decreased rate for insensitive ones. Notably, the change in privacy leakage is more pronounced than the change in privacy awareness after the incorporation of safety prompts. For example, Table VI shows that GPT-4o [OpenAI-gpt4]’s $RtA$ for sensitive questions increased by 0.14 after adding safety prompts, while the recognition rate for sensitive ones only improved by 0.032.
In the inverse enhancement pattern, we observe that incorporating safety prompts reduces the recognition rate of sensitive questions for some models, like otter [li2023ottermultimodalmodelincontext] and phi-3-vision [abdin2024phi], while increasing their recognition rate for insensitive questions. However, this inverse enhancement does not negatively impact the models’ performance in Privacy Leakage. For example, although the addition of safety prompts decreased phi-3-vision [abdin2024phi]’s recognition rate for sensitive questions by 0.20, its $RtA$ for sensitive questions actually improved by 0.18. The existence of this inverse enhancement pattern further underscores a significant disconnect between awareness and behavior in current models.
Under identical prompt engineering, Privacy Leakage and Privacy Awareness undergo divergent changes, further demonstrating that current LVLMs exhibit misalignment between their privacy awareness and privacy protection capabilities.
VI Discussion
VI-A Social and Ethical Concerns
The privacy categories encompassed by Multi-PA are formulated by a team of researchers with expertise in computer science and psychology, and offer valuable insights for evaluating the privacy preservation capabilities of Large Vision-Language Models (LVLMs). Notably, trade secret and state secret are considered highly sensitive types of confidential information. To mitigate the risk of attackers leveraging our benchmark to extract such sensitive data, we rigorously filtered the collected images, retaining only publicly available, low-sensitivity image data related to trade and nation. Specifically, we avoid collecting images that are sourced from classified websites, like WikiLeaks [^3] A platform from https://wikileaks.org/, and published within a year. We argue that these measures may help prevent our benchmark from posing significant social risks. We believe evaluating privacy risks LVLMs on these two privacy types (trade secret and state secret) is essential, as a model’s ability to protect highly sensitive data reflects its potential societal risks and demonstrates responsible consideration for its deployment and application. Multi-PA may empower model developers to gain a more granular understanding of their models’ privacy vulnerabilities and implement appropriate safety mechanisms to mitigate these risks, ultimately contributing to the development of privacy-enhanced LVLMs.
VI-B Limitation
We do not evaluate the model’s privacy preservation capabilities under privacy attack scenarios. Recent studies have demonstrated that jailbreak attacks can bypass model safety mechanisms, leading to unsafe outputs. Since jailbreak attacks do not require prior knowledge of the model’s training data and can easily be implemented through prompt engineering to extract private information, the need for privacy evaluation under jailbreak attack scenarios is both crucial and urgent. We plan to incorporate diverse jailbreak attack scenarios in subsequent iterations of our benchmark to provide a more comprehensive evaluation of models’ privacy preservation capabilities. Moreover, privacy exhibits temporal sensitivity, meaning that over time, our privacy evaluation samples may become inadequate to assess future large vision-language models. This time-dependent nature of privacy implies that evaluation samples require continuous updates as models progress. To address this, we plan to develop a dynamic and scalable privacy assessment framework capable of adapting to the advancements of LVLMs in our future work.
VII Conclusion
@@
In this paper, we introduce Multi-PA, a multi-perspective assessment on privacy evaluation of LVLMs. Based on Multi-PA, our findings indicate that GPT-4o [OpenAI-gpt4] demonstrates a promising understanding of privacy awareness, whereas Phi [abdin2024phi] exhibits superior performance in preventing privacy leakage. Further investigation into tasks in Privacy Leakage reveals heterogeneity in the models’ privacy preservation capabilities across different privacy leakage modes and privacy types (personal privacy, trade and state secrets). Moreover, stronger privacy preservation mechanisms often correlate with reduced responsiveness to privacy-unrelated queries, posing a challenge in balancing these two objectives. Multi-PA aims to systematically analyze the limitations and vulnerabilities inherent in existing privacy preservation mechanisms, and inform the development of more robust privacy-preserving models.
[Uncaptioned image] Jie Zhang (Member, IEEE) received the Ph.D. degree from the University of Chinese Academy of Sciences (CAS), Beijing, China. He is currently an Associate Professor with the Institute of Computing Technology, CAS. His research interests include computer vision, pattern recognition, machine learning, particularly include face recognition, image segmentation, weakly/semi-supervised learning, and domain generalization.
[Uncaptioned image] Xiangkui Cao (Student Member, IEEE) received the B.S. degree from University of Chinese Academy of Sciences in 2023. He is currently pursuing the Ph.D. degree from University of Chinese Academy of Sciences. His research interest includes AI safety and truthfulness.
[Uncaptioned image] Zhouyu Han received the B.S. degree in computer science and technology from the University of Chinese Academy of Sciences, China, in 2024. He is currently pursuing the master’s degree with the Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China. His research interests include security of multimodal large-language models and model fingerprints.
[Uncaptioned image] Shiguang Shan (Fellow, IEEE) received the Ph.D. degree in computer science from the Institute of Computing Technology (ICT), Chinese Academy of Sciences (CAS), Beijing, China, in 2004. He has been a Full Professor with ICT since 2010, where he is currently the Director of the Key Laboratory of Intelligent Information Processing, CAS. His research interests include signal processing, computer vision, pattern recognition, and machine learning. He has published more than 300 articles in related areas. He served as the General Co-Chair for IEEE Face and Gesture Recognition 2023, the General Co-Chair for Asian Conference on Computer Vision (ACCV) 2022, and the Area Chair of many international conferences, including CVPR, ICCV, AAAI, IJCAI, ACCV, ICPR, and FG. He was/is an Associate Editors of several journals, including IEEE Transactions on Image Processing, Neurocomputing, CVIU, and PRL. He was a recipient of the China’s State Natural Science Award in 2015 and the China’s State S&T Progress Award in 2005 for his research work.
[Uncaptioned image] Xilin Chen (Fellow, IEEE) is currently a Professor with the Institute of Computing Technology, Chinese Academy of Sciences (CAS). He has authored one book and more than 400 articles in refereed journals and proceedings in the areas of computer vision, pattern recognition, image processing, and multimodal interfaces. He is a fellow of the ACM, IAPR, and CCF. He is also an Information Sciences Editorial Board Member of Fundamental Research, an Editorial Board Member of Research, a Senior Editor of the Journal of Visual Communication and Image Representation, and an Associate Editor-in-Chief of the Chinese Journal of Computers and Chinese Journal of Pattern Recognition and Artificial Intelligence. He served as an organizing committee member for multiple conferences, including the General Co-Chair of FG 2013/FG 2018, VCIP 2022, the Program Co-Chair of ICMI 2010/FG 2024, and an Area Chair of ICCV/CVPR/ECCV/NeurIPS for more than ten times.
Supplementary Materials